OpenStudy (anonymous):

Given the function f(x) = x2 and k = –1, which of the following represents a function opening downward? f(x) + k kf(x) f(x + k) f(kx)

OpenStudy (danjs):

the function is a quadratic, parabola f(x) = a*x^2 + b*x + c here a=1, and b and c are zero

OpenStudy (danjs):

If you recall, the 'a' coefficient on the x^2 term can tell you which way the parabola opens, up+ or down -

OpenStudy (danjs):

the thing starts opening upwards f(x) = x^2 to open down, need a negative in the front

OpenStudy (danjs):

so to do that change using that k=-1 you can multiply f(x) by that k

OpenStudy (danjs):

k*f(x) = (-1)*f(x) = (-1)*x^2

OpenStudy (anonymous):

so the answer would be b?

OpenStudy (danjs):

f(x) + k --- that shifts the whole graph up/down by k, any value is just increased by k k*f(x) --- scale, also changed the opening direction depending on the sign of 'a' f(x + k) -- horizontal shifts , f(x-a) , so f(x+k) moves left k units for the whole graph

OpenStudy (anonymous):

if I’m being honest, I’m not sure what that means. What would the answer be, or could you try showing me. @DanJS

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

was gone for awhile, yeah for f(x) = x^2, which is a parabola with vertex at the origin and opening upwards, to flip it to open down, you need a negative value as a coefficient to x^2

OpenStudy (danjs):

yeah, k*f(x) = -x^2 opens down now

OpenStudy (danjs):

I just figured you probably may of went over the other ideas in the answer choices, i was typing them to show why wrong,